Phone details


Present: AK, PH, PF, MT, JT, HN, KL

  1. Actions
    • Action on CMcW to trace trigger signal to FC datalogger.
      • Stands.
    • Action MT/JB to agree and implement testing of quench valves.
      • Done. System works. Not yet clear if it has been witnesed. Action changed to ensure that test is witnessed and documented.
    • Action PW to modify purge sequence limits in agreement with JB.
      • Stands.
    • Action AK/PW to test permissions of required connections HMI to PLC.
      • Stands.
    • Action CMcW/JB to search archiver for FC radiation shield temp & helium pressure excursions.
      • Done. PH gave plot to JB. Features of quenches identified.
    • Action PF/AK to propose IOC test program for next week.
      • Done. IOC isolation has been implememted successfully. Firewall is open while we are in development mode.
    • Action AK/JB to arrange further meeting on alarm handler development
      • Done. Meeting held on 01Aug17. Parameters were specified. Spreadsheet has been updated. Outtanding issue is that parameters can not be set with firewall up. So need to hard-code limits in IOC.
    • Action JB/PF to agree wiki page for paperwork.
      • Done. This is teh document pge under operations.
    • Action JB Atex rating of everything in H2 shed.
      • Stands.
    • Action KL discuss with JT final paperwork required - note ISIS requires document giving clear instruction for ISIS MCR
      • Done. Documentation defined in the OPI provided by AN.
    • Action JB to prepare rota on 'limited call-out' basis
      • Stands.
    • Action CW to contact HR re 'on-call' times
      • Ongoing.
  2. Actions defined Preliminary H2 tour
    • Actions noted on the tour by AN:
      • Arrange final documentation review
        • Stands.
      • Completion of Hazop requirements
        • Stands.
      • Correct the labelling on absorber on GUI diagram
        • Done.
      • Apply labels to pipes in loading bay area
        • Check.
      • Secure crash barriers around charging enclosure to roadway
        • Stands.
      • Check valve panel area is correctly identified on ATEX zoning drawing and apply ATEX zone notice to area
        • Stands.
      • Check that an alarm is provided in event of extract fans/pumps failing
        • Stands.
      • Check bonding of lightning conductor
        • Stands.
      • Arrange for testing and certification of the lift-off vent
        • Stands.
    • The following actions all apply to the R5.2 LH2 roof enclosure:
      • PH is in contct with A. Gallagher and relevant contractors. Trying to push to get commitment for work to be carried out. Will cost money.
      • Arrange correct type of ATEX-compliant glands for vacuum pump motor wiring
      • Add ATEX-compliant blanks to unused ports in vacuum pump motor terminal boxes
      • Electrical connection to thermostat to be via ATEX-compliant gland
      • Fit and test earth bond connection to heaters
      • Check cable glands to gas detection heads and if necessary replace with ATEX-compliant ones
      • Verify that the gas detection heads have been calibrated for hydrogen (rather than methane) and amend calibration plates
      • Conduits for external lighting and beacons to be sealed to ATEX compliant standard
      • Electrical bonding of all metalwork and equipment should be installed and tested
      • Arrange a qualified inspection of the installation in line with the requirements of BS-EN60079-17, before the final safety tour
    • AN note on initial safety tour:
  3. Monitoring and archival of data - PF/AK
    • IOC
      • IOC now in use. Steady progress updating IOC in response to request. The only outstanding is the coding of the alarm limits.
    • Alarm handler
      • Values hvae been udated in spreadsheet, now need to migrate to alarm handler. Hard-code limits in IOC. Archiver is running.
      • Request to have heater activity to be displayed on the GUI.
  4. Neon Run
    • Progress report JB/PW/MT/HN
      • Running well, stable operation since Monday. 1.0l neon, stable at 1120mB. Next will retest empty with empty sequence.
      • Level sensors tested and they are working nicely. >100mK change when in vapour. 0mK change in liquid.
      • Will check what happens when fans are turned off.
  5. Current technical
    • Refurbishment of pipework from turret to wall.
      • P.Lester now on the job. OxFab making transfer line. EC and JT will make pipe supports. To be done by next week.
      • Will need to survey in situ to calculate length of central section of pipe.
    • Refurbishment of pipework from wall to roof and vent.
      • Need to inspect work that need to be done.
  6. Documentation
    • List compiled by AN in OPI and required at final safery tour.
      • Operational risk assessments and method statements
      • Service and maintenance manuals and/or documents
      • List of Competent Persons
      • Test certificates and statutory written schemes of inspection and declarations of conformity
      • Completed action list from the preliminary safety tour, including actions from technical reviews and HAZOP
      • Sign off sheet(s)
      • Approval of the liquid hydrogen system by the Liquid Hydrogen Working Group
  7. Timetable
  8. Next meeting
    • Thurs 03th August 2017; in the MLCR
  9. AoB

Summary of actions

  • CMcW: to trace trigger signal to FC datalogger.
  • PW: to modify purge sequence limits in agreement with JB.
  • AK/PW: to test permissions of required connections HMI to PLC.
  • MT/JB: Ensure D. Haines witnesses test of quench valve and complete appropriate documentation.
  • JB: Atex rating of everything in H2 shed.
  • JB: to prepare rota on 'limited call-out' basis
  • CW: to contact HR re 'on-call' times

Back to LH2:



Updated by Long, Kenneth almost 7 years ago · 6 revisions